OpenStudy (samigupta8):

The limiting line in balmer series will have a frequency of ?? (Rydberg constant , R=3.29*10^15 cycle/s)

OpenStudy (samigupta8):

The value of R is for infinity

OpenStudy (samigupta8):

@parthkohli

OpenStudy (mayankdevnani):

tumne jo information di h voh galat h

OpenStudy (mayankdevnani):

\[\large \bf R_{\infty}=1.09 \times 10^7\]

OpenStudy (mayankdevnani):

and the value of R you provided is \[\large \bf R_{\infty} \times c=3.29 \times 10^{15}\]

OpenStudy (mayankdevnani):

hope you now solve it :)

OpenStudy (samigupta8):

I m getting ans having order of 10^39... Bt there is no option as that

OpenStudy (samigupta8):

N even i provided you the correct info bcoz the unit of R that i gave is cycles/s.. N u r giving the constant in terms of per metre

OpenStudy (samigupta8):

Options are 3.65*10^14 8.22*10^14 3.29*10^15 5.26*10^13

OpenStudy (priyar):

its option B (since R/4)..

OpenStudy (samigupta8):

We are asked about frequency ..the formula for energy is hnu =R /4

OpenStudy (priyar):

no here they have given in cycles/sec

OpenStudy (samigupta8):

So ??

OpenStudy (priyar):

"frequency is the no. of cycles per unit time"

OpenStudy (priyar):

so we directly use R/4..ok?

OpenStudy (samigupta8):

Bt the question has not given the units in term of cycle /s Ans is in units of per s

OpenStudy (samigupta8):

@priyar

OpenStudy (priyar):

both units are correct!

OpenStudy (samigupta8):

I think they have given it wrong how can the unit of rydbergy constant be in cycle/s ..evn if we consider mayank's post the unut would be /s

OpenStudy (samigupta8):

Rydberg*

OpenStudy (priyar):

wait..i will post a bit of info. from a website..

OpenStudy (samigupta8):

Alright!!

OpenStudy (priyar):

"The SI unit of frequency is the hertz (Hz), named after the German physicist Heinrich Hertz; one hertz means that an event repeats once per second. A previous name for this unit was cycles per second (cps). The SI unit for period is the second."

OpenStudy (priyar):

Now okay?

OpenStudy (samigupta8):

I think so...

OpenStudy (mayankdevnani):

simply apply the formula :- \[\large \bf frequency=3.29 \times 10^{15}(\frac{1}{2^2}-\frac{1}{\infty^2})\]

OpenStudy (mayankdevnani):

you'll get your answer :)

OpenStudy (mayankdevnani):

i don't find any difficulty in this question !

OpenStudy (samigupta8):

I got it already @mayankdevnani

OpenStudy (mayankdevnani):

great !